OpenStudy (anonymous):

Simplify: 12√7 - 5√28 a. -8√7 b. 2 c. 2√7 d. 2√14

OpenStudy (anonymous):

c

OpenStudy (anonymous):

5√28 = 5*√4 * √7 = 10√7 so 12√7 - 5√28 = 12√7 - 10√7 = 2√7

OpenStudy (dumbcow):

\[\sqrt{28} = \sqrt{4*7} = \sqrt{4}\sqrt{7}=2\sqrt{7}\]